OpenStudy (anonymous):

how do you find the sum of 3 consecutive numbers? step by step plz the sum of 3 consecutive numbers is 183. find the numbers.

OpenStudy (mathteacher1729):

Do you know what the word "consecutive" means?

OpenStudy (saifoo.khan):

First = x Second = x+1 third = x+2 therefore. x+ ( x+1) + (x+2) = 183. Solve.

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

i think you are not looking to find the sum of three consecutive numbers, but three consecutive numbers whose sum is 183

OpenStudy (anonymous):

ohhh okay , so all you need to do it find 3 numbers that equal 183?

OpenStudy (anonymous):

x+x+1+x+2 = 183 3x + 3 = 183 3x = 183 - 3 = 180 x = 180/3 = 60

OpenStudy (anonymous):

so no.s are 60 , 61 & 62